OpenStudy (anonymous):

x3 – x2 + 5x – 5 = 0 (1 point)

OpenStudy (anonymous):

Answer Choices i sqrt 5, -i sqrt 5, 1 i sqrt 5, -i sqrt 5, -1 sqrt 5, - sqrt 5, -1 i sqrt 5, 1

hero (hero):

x^3 - x^2 + 5x - 5 = 0 x^2(x - 1) + 5(x - 1) = 0 (x - 1)(x^2 + 5) = 0 x - 1 = 0 x^2 + 5 = 0 x = 1 x^2 = -5 x = ± √(-5) x = ± √(5)i

OpenStudy (anonymous):

Thank You ! :D